OpenStudy (anonymous):

Find the volume of a prism if the length is 2p-3 the width is p+2 and the height is p-4

Directrix (directrix):

Is this a rectangular prism?

Directrix (directrix):

@volleyballgirl13

OpenStudy (campbell_st):

volume is \[V = length.. \times.. width.. \times height..\] so \[V = (2p -3) \times (p + 2) \times (p - 4)\] just multiply it out for a cubic equation.

OpenStudy (anonymous):

Yes a rectangular prism.

Directrix (directrix):

@VolleyballGirl13 Do you have an idea as to what the answer might be, assuming you did the multiplication above?

OpenStudy (anonymous):

I got 2p³-5

Directrix (directrix):

What did you get when you multiplied (2p - 3) times (p + 2) ?

Directrix (directrix):

I think I see your error. (2p-3) (p+2) (p-4) You, I think, multiplied 2p times p times p to get 2p^3. Then, you added -3 and 2 and -4. That is not how binomials are correctly multiplied.

OpenStudy (campbell_st):

just do it in parts.. look at the 1st 2 binomials \[(2p - 3)(p + 2) = 2p^2 + p - 6\] so you need to distribute \[(2p^2 + p - 6) \times (p - 4)\]